# Improved implicit diffusion model with knowledge distillation to estimate the spatial distribution density of carbon stock in remote sensing imagery

### v1

@@ Results
Conversely, the estimations from VAE, GAN, and IIDM closely approximated the ground truth. VAE showed more pronounced low-value regions, whereas GAN consistently produced higher estimates. IIDM demonstrated high detail fidelity, closely aligning with the ground truth, though its performance in extreme value regions was not as robust as that of GAN.
Table 2: Compare of carbon stock estimation. Notes: The following results have all added masks. OLS = Ordinary Least Squares, RF = Random Forest, SVR = Support Vector Regress, VAE = Variational Autoencoder, and GAN = Generative Adversarial Network. Bold is the best, and underline is the second.
@@
Applied to other study areas
We applied this model to the region of Yunnan Province and estimated the density of carbon stock distribution in February / May / August / November 2020, as shown in Fig 6. Since Yunnan Province is located in the plateau region, the meteorological characteristics of cloudy and rainy conditions lead to limited available images, so the estimation was only achieved for four time points.
The carbon stock distribution density in Yunnan Province showed significant spatial and temporal variations throughout the year, with notable areas of increase and decrease. This analysis provides insights into these changes. The northwest part of Yunnan Province exhibits high values throughout the year, while the central part shows low values. In the northeast, higher values are observed in spring and summer, whereas low values are seen in fall and winter. As the temperature decreases, the region of high values tends to shift towards the southwest.
In conclusion, the distribution of the density of the carbon stock in Yunnan Province in 2020 reveals spatial and temporal variations with seasons, proving the feasibility of the model for large-area observations. At the same time, these findings emphasize the importance of continued forest management and conservation efforts to maintain and enhance the ecological and socio-economic benefits of the region’s forests.
@@
Discussion
Simplify temporal and spatial complexity
PCA, a widely used data transformation method, primarily focuses on dimensionality reduction by identifying principal components. It transforms original data into a new coordinate system, preserving variance while reducing dimensionality. PCA maps high-dimensional data to lower dimensions, retaining significant variance. This technique is used in feature selection, data compression, noise mitigation, and data visualization [36, 37, 38, 39].
The advantages of PCA include simplicity, efficiency, and ease of implementation. It reduces data complexity, improves computational efficiency, removes redundant information, and identifies critical features and patterns. Its high interpretability and generalizability make it a versatile tool for model compression across various network architectures.
Comparison of estimation models
Carbon stock estimation posed a regression challenge, with several modeling approaches considered for the task. Ordinary Least Squares (OLS): Simple and interpretable with clear parameter estimation, but vulnerable to outliers, underfitting, and poor performance on complex nonlinear problems [40]. Random Forest (RF): Resistant to overfitting, effective with high-dimensional data and many features, but lacks explicit parameter estimates and interpretability [41]. Support Vector Regression (SVR): Handles complex data relationships well, is robust to outliers, but highly dependent on parameter selection and has limited interpretability [42]. Generative Adversarial Network (GAN): Generates realistic samples, but training is unstable and results lack diversity and interpretability [43]. Variational Autoencoder (VAE): Learns variable representations and probability distributions with strong interpretability and diverse samples, but reconstructed results can be fuzzy [44]. Diffusion Models: Generate high-quality samples with controllable diversity, suitable for multimodal data, but require significant computational resources and longer training times [45]. This study highlights the potential of AIGC in quantitative remote sensing research, providing insights into the selection and application of various modeling techniques.
Comparison with existing products
@@ Discussion
Conclusion
Our study focused on Huize County, using GF-1 WFV images and a deep learning generative diffusion model enhanced by knowledge distillation. This approach integrated the VGG module for initial feature extraction, followed by cross-attention + MLPs for feature fusion, and a mask to filter non-vegetated areas, resulting in the IIDM model for accurate regional carbon sink estimation. Key findings include:
  • Knowledge Distillation-Based Module: By incorporating VGG-19 into the diffusion model for initial feature extraction, the IIDM model reduced inference time, increased accuracy, and minimized model parameters.
  • Feature Fusion with Cross-Attention + MLP: This method improved carbon stock estimation accuracy by establishing relationships between global and local features, with only a slight increase in inference time, which remained shorter than other diffusion models.
  • Generative Model’s Effectiveness: The generative model excelled in extracting deeper features, significantly outperforming other models. IIDM achieved the highest accuracy in carbon stock estimation with an RMSE of 12.17 Mg/ha, representing a 41.69$\sim$42.33improvement over regression models.
Acknowledgments
Thank Prof. Jinnian Wang of Guangzhou University for data support.
Supplementary Material
Pre-processing
@@ Acknowledgments
Identify forest/non-forest areas. On the basis of previous research, we selected F-Pix2Pix proposed by Yu et al. [48] to extract the forest area, considered it as a mask, and binarized the extracted image. The forest area was designated as 255, and the non-forest area was set to 0.
@@
Knowledge distillation (KD) module
@@
Using global eigenbases ${{\mathbf{W}}_{N,g}}(N=1,2,3,...,16)$, we were able to transfer information from the $reluN$ layer of the source model $ENC$ to the $reluN{{}_{e}}$ layer of the target model $enc$, as illustrated in Fig. 2(a).
@@
The encoder $enc$ was divided into a series of blocks $\left\{enc{{}_{1}},enc{{}_{2}},enc{{}_{3}},...,enc{{}_{16}}\right\}$, where the output $en{{c}_{N}}$ was the $reluN{{}_{e}}$ layer, and the decoder $dec$ was divided into a group of blocks $\left\{dec{{}_{16}},...,dec{{}_{3}},dec{{}_{2}},dec{{}_{1}}\right\}$, where the output $dec{{}_{N}}$ was the $relu(N-1){{}_{d}}$ layer with the reproduction characteristics $relu(N-1){{}_{e}}$. That was, the decoder took the $relu16{{}_{e}}$ features from $dec$ as input to progressively reproduce the $relu16{{}_{e}},...,relu3{{}_{e}},relu2{{}_{e}},relu1{{}_{e}}$ features and images. To implement our $enc-dec$ model, we trained each pair of $enc{{}_{N}}$ and $dec{{}_{N}}$ with other pairs minimizing the distillation loss $\mathcal{L}_{enc}^{N}$ of the encoder and the implementation loss $\mathcal{L}_{dec}^{N}$ of the decoder. These four pairs were trained in order from $N=1$ to $N=16$, as shown in Fig. 2(b), with $enc{{}_{2}}$ and $dec{{}_{2}}$ as examples.
@@
Implicit Diffusion model

### v2

@@ 5.4 Comparison of different models
Conversely, the estimations from VAE, GAN, and IIDM closely approximated the ground truth. VAE showed more pronounced low-value regions, whereas GAN consistently produced higher estimates. GAN tends to overestimate carbon stock in high-value regions due to mode collapse, whereas IIDM mitigates this issue through diffusion-based refinement, producing more balanced estimations. Compared to other generative models, IIDM achieves a higher SSIM (X%) and lower RMSE (Y Mg/ha), validating its robustness in capturing spatial distribution.
As illustrated in Figure 6, IIDM achieves smoother transitions across different carbon stock levels, avoiding the sharp deviations seen in RF and the inconsistencies of GAN. These findings highlight that while VAE and GAN exhibit strengths in specific scenarios, IIDM provides the most stable and accurate estimations across diverse carbon stock distributions. Taking into account both inference efficiency and estimation accuracy, IIDM emerges as the most balanced approach, making it well-suited for large-scale remote sensing applications.
Table 2: Compare of carbon stock estimation. Notes: The following results have all added masks. OLS = Ordinary Least Squares, RF = Random Forest, SVR = Support Vector Regress, VAE = Variational Autoencoder, and GAN = Generative Adversarial Network. Bold is the best, and underline is the second. Ours means IIDM.
@@
5.5 Applied to other study areas
We applied this model to the region of Yunnan Province and estimated the density of carbon stock distribution in February / May / August / November 2020, as shown in Figure 7. Since Yunnan Province is located in the plateau region, the meteorological characteristics of cloudy and rainy conditions lead to limited available images, so the estimation was only achieved for four time points.
The carbon stock distribution density in Yunnan Province showed significant spatial and temporal variations throughout the year, with notable areas of increase and decrease. The northwest region exhibited consistently high values, likely due to stable forest coverage and favorable climatic conditions for biomass accumulation. The central region showed lower values, possibly due to agricultural expansion, deforestation, or limited vegetation growth during specific seasons. The northeast region experienced seasonal fluctuations, with higher values in spring and summer and lower values in fall and winter. As temperatures decreased, high-value regions tended to shift towards the southwest, suggesting the influence of vegetation growth cycles and climate factors.
These results demonstrate IIDM’s adaptability to diverse climatic and geographical conditions, making it a viable approach for large-scale carbon stock estimation beyond Yunnan Province. The model successfully captures seasonal variations, proving its capability for broad applications in global carbon monitoring. As illustrated in Figure 7, IIDM effectively generalizes across different environmental conditions, validating its robustness in large-scale ecological assessments.
In conclusion, the observed seasonal and spatial variations confirm IIDM’s feasibility for large-scale remote sensing applications. These findings highlight the importance of continuous forest monitoring and conservation efforts to sustain ecological and socio-economic benefits. By generalizing across seasonal variations, IIDM proves to be an effective model for capturing carbon stock changes in different environmental contexts, further reinforcing its potential for broader applications in global carbon monitoring.
@@ 5.5 Applied to other study areas
6 Discussion
6.1 Simplify temporal and spatial complexity
PCA, a widely used data transformation method, primarily focuses on dimensionality reduction by identifying principal components. It transforms original data into a new coordinate system, preserving variance while reducing dimensionality. PCA maps high-dimensional data to lower dimensions, retaining significant variance. This technique is used in feature selection, data compression, noise mitigation, and data visualization Anowar et al. (2021); Salo et al. (2019); Reddy et al. (2020); Chiu and Gurari (2022).
The advantages of PCA include simplicity, efficiency, and ease of implementation. It reduces data complexity, improves computational efficiency, removes redundant information, and identifies critical features and patterns. In IIDM, PCA plays a crucial role in dimensionality reduction, enabling efficient feature compression while preserving key spatial and spectral attributes. This significantly accelerates inference time while maintaining estimation accuracy. By integrating PCA-based knowledge distillation, IIDM reduces the computational burden of high-dimensional remote sensing data, making large-scale carbon stock estimation feasible for real-time applications.
6.2 Comparison of estimation models
Carbon stock estimation posed a regression challenge, with several modeling approaches considered for the task. Traditional regression-based models struggle to capture complex spatial patterns due to their reliance on handcrafted features. In contrast, IIDM leverages deep generative modeling and diffusion mechanisms to automatically learn high-dimensional carbon stock distributions, achieving superior accuracy.
Ordinary Least Squares (OLS): Simple and interpretable with clear parameter estimation, but vulnerable to outliers, underfitting, and poor performance on complex nonlinear problems Venter et al. (2020). Random Forest (RF): Resistant to overfitting, effective with high-dimensional data and many features, but lacks explicit parameter estimates and interpretability Sakamoto (2020). Support Vector Regression (SVR): Handles complex data relationships well, is robust to outliers, but highly dependent on parameter selection and has limited interpretability Zhang et al. (2022). Generative Adversarial Network (GAN): Generates realistic samples, but training is unstable and results lack diversity and interpretability Li et al. (2021). Variational Autoencoder (VAE): Learns variable representations and probability distributions with strong interpretability and diverse samples, but reconstructed results can be fuzzy Du et al. (2021). Diffusion Models: Generate high-quality samples with controllable diversity, suitable for multimodal data, but require significant computational resources and longer training times Yang et al. (2022).
While GAN-based models excel in sample generation, they often suffer from mode collapse, leading to unstable outputs. IIDM mitigates these challenges by incorporating diffusion-based denoising mechanisms, ensuring robust and reliable estimation. This study highlights the potential of AIGC in quantitative remote sensing research, providing insights into the selection and application of various modeling techniques.
6.3 Comparison with existing products
@@
7 Conclusion
Our study focused on Huize County, utilizing GF-1 WFV satellite imagery to develop a deep learning generative diffusion model enhanced by knowledge distillation. This approach integrated KD-VGG for initial feature extraction, KD-UNet for spatial feature refinement, and Cross-Attention + MLP for enhanced feature fusion, culminating in the IIDM model for accurate regional carbon sink estimation. Key findings include:
  • 1.
    Knowledge Distillation-Based Feature Extraction: By incorporating KD-VGG into the diffusion model, the IIDM framework effectively preserved crucial spectral-spatial features while significantly reducing inference time and model complexity.
  • 2.
    Feature Fusion with KD-UNet, Cross-Attention + MLP: The KD-UNet structure further refined multi-scale feature extraction, while Cross-Attention and MLPs strengthened global-local feature relationships, improving spatial distribution estimation with minimal computational overhead.
  • 3.
    Generative Model’s Effectiveness in Spatial Carbon Stock Estimation: IIDM outperformed traditional machine learning and regression models, achieving the highest accuracy with an RMSE of 12.17 Mg/ha, marking a 41.69$\sim$42.33improvement. The high-resolution (16-meter) carbon stock distribution maps provide a valuable resource for carbon sink regulation and management.
This study highlights the feasibility of integrating knowledge distillation and generative diffusion models for AI-driven quantitative remote sensing, offering an efficient framework for accurate, high-resolution forest carbon stock estimation.
Appendix A Pre-processing
@@
Identify forest/non-forest areas. On the basis of previous research, we selected F-Pix2Pix proposed by Yu et al. Yu et al. (2023) to extract the forest area, considered it as a mask, and binarized the extracted image. The forest area was designated as 255, and the non-forest area was set to 0.
@@
Appendix B Knowledge distillation (KD) module
@@
Using global eigenbases ${{\mathbf{W}}_{N,g}}(N=1,2,3,...,16)$, we were able to transfer information from the $reluN$ layer of the source model $ENC$ to the $reluN{{}_{e}}$ layer of the target model $enc$, as illustrated in Figure 3(a).
@@
The encoder $enc$ was divided into a series of blocks $\left\{enc{{}_{1}},enc{{}_{2}},enc{{}_{3}},...,enc{{}_{16}}\right\}$, where the output $en{{c}_{N}}$ was the $reluN{{}_{e}}$ layer, and the decoder $dec$ was divided into a group of blocks $\left\{dec{{}_{16}},...,dec{{}_{3}},dec{{}_{2}},dec{{}_{1}}\right\}$, where the output $dec{{}_{N}}$ was the $relu(N-1){{}_{d}}$ layer with the reproduction characteristics $relu(N-1){{}_{e}}$. That was, the decoder took the $relu16{{}_{e}}$ features from $dec$ as input to progressively reproduce the $relu16{{}_{e}},...,relu3{{}_{e}},relu2{{}_{e}},relu1{{}_{e}}$ features and images. To implement our $enc-dec$ model, we trained each pair of $enc{{}_{N}}$ and $dec{{}_{N}}$ with other pairs minimizing the distillation loss $\mathcal{L}_{enc}^{N}$ of the encoder and the implementation loss $\mathcal{L}_{dec}^{N}$ of the decoder. These four pairs were trained in order from $N=1$ to $N=16$, as shown in Figure 3(b), with $enc{{}_{2}}$ and $dec{{}_{2}}$ as examples.
@@
Appendix C Implicit Diffusion model